H0:p=0.63 H1:p>0.63 Your sample consists of 100 subjects, with 67 successes. Calculate the test statistic, rounded to 2 decimal places z=
Answer #1

solution

р = x / n = 67 / 100=0.67

P0 = 0.63

1 - P0 = 1 - 0.63=0.37

Test statistic = z

= р - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.67 - 0.63/ [\sqrt(0.63*0.37) /100 ]

z = 0.83